The neighborhood of x,if |2x+9|<25,is


1
Expert's answer
2021-04-28T05:15:21-0400

 |2x+9|<25

=> "-25<2x+9<25"

Adding -9 we get

"-25-9<2x+9-9<25-9"

=> "-34<2x<16"

Dividing by 2

"\frac{-34}{2}<\frac{2x}{2}<\frac{16}{2}"

=> "-17<x<8"

So the neighborhood of x is "(-17, 8)"